OpenStudy (anonymous):

write the equation of the line that passes through ( -2 , 1 ) and ( 5 , 4 ) any help ????

OpenStudy (anonymous):

first find the slop of the line

OpenStudy (anonymous):

ok , and then after that ? thanks for the help btw

OpenStudy (anonymous):

easy part. we know that y=mx+b right ? so then we can find the y-intercept and the x-intercept .

OpenStudy (anonymous):

so then , that would be the equation ?

OpenStudy (anonymous):

try it and see.

OpenStudy (anonymous):



OpenStudy (anonymous):

I got that the equation turned to be y= 3/7x + 13/7 because I found the slope and then did y= mx + b ... did I do it right ?!?!

OpenStudy (anonymous):

thats right :)

OpenStudy (anonymous):

thank you very much :)